solve the equation in exact differentials

(2-9xy2)xdx+(4y2-6x3)ydy=0


1
Expert's answer
2022-02-08T07:25:52-0500

Let us solve the equation in exact differentials "(2-9xy^2)xdx+(4y^2-6x^3)ydy=0."


It follows that there exists the function "u=u(x,y)" such that


"\frac{\partial u}{\partial x}=(2-9xy^2)x=2x-9x^2y^2, \ \frac{\partial u}{\partial y}=(4y^2-6x^3)y=4y^3-6x^3y."


Therefore, "u=x^2-3x^3y^2+c(y)," and hence


"\frac{\partial u}{\partial y}=-6x^3y+c'(y)=4y^3-6x^3y."


It follows that "c'(y)=4y^3," and thus "c(y)=y^4+C."


We conclude that the general solution of the differential equation is


"x^2-3x^3y^2+y^4=C."